A rectangular yard has an area of 2160 square feet. The length of the yard is three feet longer than its width. What is the perimeter of the yard?

Answer by checkley79(3341)

AREA=L*W
L=W+3
2,160=(W+3)W
2,160=W^2+3W
W^2+3W-2,160=0
(W+48)(W-45)=0
W-45=0
W=45 ANS. FOR THWE WIDTH.
L=45+3=48 ANS. FOR THE LENGTH.
PROOF:
2,160=45*48
2,160=2,160
P=2L+2W
P=2*48+2*45
P=96+90
P=186 ANS. FOR THE PERIMETER.